# FALCON: Fine-grained Activation Manipulation by Contrastive Orthogonal Unalignment for Large Language Model

### v1

@@ 6 Conclusion
This paper presents FALCON, a fine-grained representation-guided framework for LLM unlearning. Leveraging mutual information guidance and contrastive orthogonal unalignment, it enables precise and efficient unlearning through principal component-based representation separation and gradient conflict resolution. Extensive experiments demonstrate its superior performance in effectively removing undesired knowledge while preserving essential information, along with resistance against knowledge recovery and efficient optimization guidance. However, this work is currently limited to text-based LLM unlearning, with experiments conducted on relatively smaller models due to computational constraints. Future directions include extending unlearning to multi-modal LLMs and refining strategies to disentangle intertwined knowledge in deeper architectures.
@@ Appendix A Information-Theoretic Metrics
Information-theoretic metrics, such as mutual information and entropy, provide a robust theoretical foundation for understanding and managing uncertainty in machine learning models, including LLMs (Malinin & Gales, 2018; Dombrowski & Corlouer, 2024). Entropy, as a measure of uncertainty, has been widely applied to assess prediction informativeness (Van Amersfoort et al., 2020), guide feature selection (Deng et al., 2022), and reduce predictive uncertainty (Malinin & Gales, 2018). In LLMs, entropy-based methods have been used to evaluate model confidence, regularize outputs, and detect hallucinations (Attanasio et al., 2022; Qiu & Miikkulainen, 2024; Farquhar et al., 2024). Similarly, mutual information quantifies shared information between variables, offering a principled approach to analyzing dependencies within model layers, improving representation learning, and understanding information propagation across deep neural networks (Gabrié et al., 2018; Tschannen et al., 2020). In LLMs, MI has been leveraged to optimize pretraining objectives, identify task-relevant variables during fine-tuning, and improve knowledge distillation (Cha et al., 2022; Wu et al., 2024; Chen et al., 2024). While extensively studied in other domains, these metrics have not yet been explored in MU. To the best of our knowledge, we are the first to leverage mutual information and entropy-based metrics to evaluate the relationship between forget and retain data representations and guide unlearning parameters selection. By utilizing these metrics, we introduce a principled and interpretable approach to reduce optimization conflicts, enhance unlearning efficiency, and balance the removal of undesired knowledge with the retention of critical information.
@@ Appendix B Contrastive Learning & Gradient Projection
Contrastive learning has emerged as a key technique for representation learning, leveraging the principle of maximizing similarity between positive pairs while minimizing it for negative pairs (Hu et al., 2024). It has shown success in self-supervised learning, feature disentanglement, and robustness improvement in deep neural networks (Ericsson et al., 2022; Kahana & Hoshen, 2022; Wang et al., 2024). Recent works have explored its application in MU, where it is used to suppress target representations while preserving critical functionality (Kim & Woo, 2022b; Yang & Li, 2023). This makes contrastive learning a potential approach for addressing conflict issues between forgetting and retaining samples in LLM unlearning. Gradient projection, on the other hand, addresses optimization conflicts by projecting gradients onto feasible directions aligned with Pareto-optimal solutions (Iskander et al., 2023). It has been successfully applied to multi-objective tasks and continual learning, effectively achieving gradient equilibrium and ensuring stable updates (Chen et al., 2022; Zhang et al., 2024c). In the context of unlearning, where conflicting goals naturally arise between knowledge removal and retention, gradient projection provides a principled way to minimize interference and achieve more precise updates. Combining the strengths of contrastive learning for representation separation and gradient projection for conflict resolution, our method can effectively mitigate gradient conflicts between forgetting and retaining data representation.
@@ C.1 Continuous Entropy
The concept of entropy in the continuous setting, often referred to as differential entropy, measures the uncertainty of a continuous random variable (Garbaczewski, 2006; Yeung & Yeung, 2008). For a random variable $\mathcal{F}$ with probability density function $p(\mathcal{F})$, the entropy $H(\mathcal{F})$ is defined as:
[…]
@@ C.2 Joint Entropy
[…]
@@ Appendix D Implementation Details
D.1 LLMU
@@
Following RMU (Li et al., 2024), we made several modifications to LLMU (Yao et al., 2023) to better align it with our tasks. Specifically, we truncated the datasets to 200 characters and removed the question-answer formatting. Additionally, we trained LLMU using LoRA (Hu et al., 2022) with a rank of 32 and a scaling factor of 16. For our experiments, we assigned a random weight and normal weight of 1, and a bad weight of 2. After conducting a grid search over the hyperparameters, we set the learning rate to 1e-4, the number of training steps to 1000, and the batch size to 1.
D.2 SCRUB
@@
We adapted the Scalable Remembering and Unlearning unBound (SCRUB) (Kurmanji et al., 2024) framework to align with our tasks. Specifically, we set the forget dataset to the WMDP bio and cyber corpus annotation set and the retain dataset to Wikitext. SCRUB was trained using the Adam optimizer with a weight decay of 0.01 and a learning rate of 1e-4. We employed log perplexity on Wikitext as the task-specific loss. Besides, to balance the loss weightings between knowledge distillation and the task-specific loss, we tuned the $\alpha$ hyperparameter with values $[{1\times 10^{-4},1\times 10^{-3},1\times 10^{-2},1\times 10^{-1},1,10}]$.
D.3 SSD
@@
We adapted the Selective Synaptic Dampening  (Foster et al., 2024) method to make it suitable for large language models. Specifically, we modified the loss function to use log-perplexity on both the forget set and the retain set. Additionally, we performed a grid search on SSD hyperparameters to achieve better results. The grid search included thresholds of [0.1,0.5,1.0,5.0] and dampening constants of $[{1\times 10^{-4},1\times 10^{-3},1\times 10^{-2},1\times 10^{-1},1}]$.
D.4 RMU
@@
For RMU implementation, our parameter selection was followed by both Li et al.’s empirical findings (Li et al., 2024) and our mutual information visualization results, which consistently indicated layer $l=7$ as optimal for minimizing parameter entanglement. Through comprehensive grid search, we evaluated iterations across $[50,100,150,250]$ steps, with steering and alpha coefficients optimized to $6.5$ and $1150$ for Zephyr-7B, and $40$ and $200$ for Yi-6B respectively. Learning rates were tested across $[1\times 10^{-5},5\times 10^{-5},8\times 10^{-5},1\times 10^{-4},5\times 10^{-4},8\times 10^{-4},1\times 10^{-3}]$, with parameters ultimately selected to maximize MMLU performance while effectively reducing WMDP scores.
D.5 FALCON
@@
For FALCON’s implementation, we maintained comparable learning rate ranges and number of iterations to RMU. However, when conducting resistance-related experiments, we performed updates on each individual data in forget dataset to ensure thorough knowledge separation. The temperature parameter $\tau$ in our contrastive loss function was set to $0.7$. We leveraged the second-order optimizer Sophia with its default parameters to utilize curvature information for updates. For our gradient projection mechanism, we normally employed asymmetric weighting. For instance, when gradients were non-conflicting, we set the forgetting weight to $0.8$ and retention weight to $1.2$; in cases of gradient conflict, these values were adjusted to $0.2$ and $1.8$ respectively. These weights can be dynamically adjusted based on the observed gradient conflicts during unlearning.
@@ E.2 Performance Breakdown Analysis of MMLU and WMDP
Humanities $56.85\pm 0.68$ $55.86\pm 0.68$
   Formal Logic $45.24\pm 4.45$ $444.44\pm 4$ 4.44
@@
E.3 Evaluation of Recovery Resistance in Chat Settings
@@
E.4 Example of Jailbreaking Prefixes Generated by Enhanced GCG

### v2

@@ 7 Conclusion
This paper presents FALCON, a fine-grained representation-guided framework for LLM unlearning. Leveraging mutual information guidance and contrastive orthogonal unalignment, it enables precise and efficient unlearning through principal component-based representation separation and gradient conflict resolution. Extensive experiments demonstrate its superior performance in effectively removing undesired knowledge while preserving essential information across diverse tasks, along with resistance against knowledge recovery and efficient optimization guidance. However, this work is currently limited to text-based LLM unlearning, with experiments conducted on relatively smaller models due to computational constraints. Future directions include extending unlearning to multi-modal LLMs and refining strategies to disentangle intertwined knowledge in deeper architectures.
Acknowledgment
This work is partially funded by the European Union (under grant agreement ID 101212818). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or European Health and Digital Executive Agency (HADEA). Neither the European Union nor the granting authority can be held responsible for them. This work is partially supported by Innovate UK through AI-PASSPORT under Grant 10126404. Yi’s contribution is partially supported through the Royal Society international exchanges programme and in part by the Engineering and Physical Sciences Research Council [grant number EP/Y009800/1], through funding from RAi UK.
@@ Appendix A Information-Theoretic Metrics
Information-theoretic metrics, such as mutual information and entropy, provide a robust theoretical foundation for understanding and managing uncertainty in machine learning models, including LLMs [57, 12]. Entropy, as a measure of uncertainty, has been widely applied to assess prediction informativeness [78], guide feature selection [11], and reduce predictive uncertainty [57]. In LLMs, entropy-based methods have been used to evaluate model confidence, regularize outputs, and detect hallucinations [4, 64, 21]. Similarly, mutual information quantifies shared information between variables, offering a principled approach to analyzing dependencies within model layers, improving representation learning, and understanding information propagation across deep neural networks [23, 74]. In LLMs, MI has been leveraged to optimize pretraining objectives, identify task-relevant variables during fine-tuning, and improve knowledge distillation [7, 81, 10]. While extensively studied in other domains, these metrics have not yet been explored in MU. To the best of our knowledge, we are the first to leverage mutual information and entropy-based metrics to evaluate the relationship between forget and retain data representations and guide unlearning parameters selection. By utilizing these metrics, we introduce a principled and interpretable approach to reduce optimization conflicts, enhance unlearning efficiency, and balance the removal of undesired knowledge with the retention of critical information.
@@ Appendix B Contrastive Learning & Gradient Projection
Contrastive learning has emerged as a key technique for representation learning, leveraging the principle of maximizing similarity between positive pairs while minimizing it for negative pairs [30]. It has shown success in self-supervised learning, feature disentanglement, and robustness improvement in deep neural networks [19, 45, 80]. Recent works have explored its application in MU, where it is used to suppress target representations while preserving critical functionality [46, 83, 90]. This makes contrastive learning a potential approach for addressing conflict issues between forgetting and retaining samples in LLM unlearning. Gradient projection, on the other hand, addresses optimization conflicts by projecting gradients onto feasible directions aligned with Pareto-optimal solutions [37]. It has been successfully applied to multi-objective tasks and continual learning, effectively achieving gradient equilibrium and ensuring stable updates [9, 92]. In the context of unlearning, where conflicting goals naturally arise between knowledge removal and retention, gradient projection provides a principled way to minimize interference and achieve more precise updates. Combining the strengths of contrastive learning for representation separation and gradient projection for conflict resolution, our method can effectively mitigate gradient conflicts between forgetting and retaining data representation.
@@ C.1 Continuous Entropy
The concept of entropy in the continuous setting, often referred to as differential entropy, measures the uncertainty of a continuous random variable [24]. For a random variable $\mathcal{F}$ with probability density function $p(\mathcal{F})$, the entropy $H(\mathcal{F})$ is defined as:
$H(\mathcal{F})=-\int p(\mathcal{F})\log p(\mathcal{F})d\mathcal{F}$ (15)
@@ C.2 Joint Entropy
$H(\mathcal{F},\mathcal{R})=-\int\int p(\mathcal{F},\mathcal{R})\log p(\mathcal{F},\mathcal{R})\,d\mathcal{F}\,d\mathcal{R}$ (16)
@@ Appendix D Implementation Details
D.1 Algorithm Overview
We summarize the core workflow of FALCON in Algorithm 1. The algorithm aims to selectively remove unwanted knowledge from large language models by guiding updates toward task-relevant, disentangled directions. It begins by identifying candidate intervention parameters with minimal knowledge entanglement between the forget and retain sets using mutual information estimates. Once the most suitable parameters are selected, FALCON applies contrastive representation unlearning via principal offset vectors to steer activations away from undesired components, followed by orthogonalizing gradient to resolve optimization conflicts between forgetting and retention objectives. The algorithm proceeds iteratively, updating only a subset of parameters to achieve efficient and robust unlearning without requiring full retraining or access to the original dataset.
Algorithm 1 Fine-grained Activation Manipulation by Contrastive Orthogonal Unalignment
1:Pretrained model $\mathcal{M}$ with parameters $\theta$
2:Forget set $\mathcal{D}_{\mathcal{F}}$, retain set $\mathcal{D}_{\mathcal{R}}$
3:Top-K components, unlearning steps $T$, loss weights $\alpha,\beta$
4:Updated model $\mathcal{M}^{\prime}$ with target knowledge forgotten
5:for each candidate layer do
6:  Extract activations from $\mathcal{D}_{\mathcal{F}}$ and $\mathcal{D}_{\mathcal{R}}$
7:  Estimate mutual information between them
8:end for
9:Select the parameters with lowest mutual information $\triangleright$ Identified once per LLM and held fixed during unlearning
10:Extract activations at selected layer for $\mathcal{D}_{\mathcal{F}}$
11:Obtain top-K directions of principal components
12:Construct POVs to steer model activations away from dominant principal subspaces associated with undesired knowledge.
13:for step = 1 to $T$ do
14:  Sample minibatch $\mathcal{B}_{\mathcal{F}}\sim\mathcal{D}_{\mathcal{F}}$, $\mathcal{B}_{\mathcal{R}}\sim\mathcal{D}_{\mathcal{R}}$
15:  Compute contrastive loss $\mathcal{L}_{\mathcal{F}}$ and gradient $\nabla\mathcal{L}_{\mathcal{F}}$ from $\mathcal{B}_{\mathcal{F}}$
16:  Compute retention loss $\mathcal{L}_{\mathcal{R}}$ and gradient $\nabla\mathcal{L}_{\mathcal{R}}$ from $\mathcal{B}_{\mathcal{R}}$
17:  if gradients conflict then
18:   Project $\nabla\mathcal{L}_{\mathcal{F}}$ onto subspace orthogonal to $\nabla\mathcal{L}_{\mathcal{R}}$
19:  end if
20:  Combine gradients: $\nabla\mathcal{L}=\alpha\cdot\nabla\mathcal{L}_{\mathcal{F}}+\beta\cdot\nabla\mathcal{L}_{\mathcal{R}}$
21:  Update $\theta\leftarrow\theta-\eta\cdot\nabla\mathcal{L}$
22:end for
23:return Updated model $\mathcal{M}^{\prime}$
D.2 Harmful Knowledge Unlearning
D.2.1 LLMU
@@
Following RMU [50], we made several modifications to LLMU [84] to better align it with our tasks. Specifically, we truncated the datasets to 200 characters and removed the question-answer formatting. Additionally, we trained LLMU using LoRA [29] with a rank of 32 and a scaling factor of 16. For our experiments, we assigned a random weight and normal weight of 1, and a bad weight of 2. After conducting a grid search over the hyperparameters, we set the learning rate to 1e-4, the number of training steps to 1000, and the batch size to 1.
D.2.2 SCRUB
@@
We adapted the Scalable Remembering and Unlearning unBound (SCRUB) [48] framework to align with our tasks. Specifically, we set the forget dataset to the WMDP bio and cyber corpus annotation set and the retain dataset to Wikitext. SCRUB was trained using the Adam optimizer with a weight decay of 0.01 and a learning rate of 1e-4. We employed log perplexity on Wikitext as the task-specific loss. Besides, to balance the loss weightings between knowledge distillation and the task-specific loss, we tuned the $\alpha$ hyperparameter with values $1\times 10^{-4},1\times 10^{-3},1\times 10^{-2},1\times 10^{-1},1,10$}].
D.2.3 SSD
@@
We adapted the Selective Synaptic Dampening  [22] method to make it suitable for large language models. Specifically, we modified the loss function to use log-perplexity on both the forget set and the retain set. Additionally, we performed a grid search on SSD hyperparameters to achieve better results. The grid search included thresholds of [0.1,0.5,1.0,5.0] and dampening constants of $1\times 10^{-4},1\times 10^{-3},1\times 10^{-2},1\times 10^{-1},1$}].
D.2.4 RMU
@@
For RMU implementation, our parameter selection was followed by both Li et al.’s empirical findings [50] and our mutual information visualization results, which consistently indicated layer $l=7$ as optimal for minimizing parameter entanglement. Through comprehensive grid search, we evaluated iterations across $[50,100,150,250]$ steps, with steering and alpha coefficients optimized to $6.5$ and $1150$ for Zephyr-7B, and $40$ and $200$ for Yi-6B respectively. Learning rates were tested across $1\times 10^{-5},5\times 10^{-5},8\times 10^{-5},1\times 10^{-4},5\times 10^{-4},8\times 10^{-4},1\times 10^{-3}]$, with parameters ultimately selected to maximize MMLU performance while effectively reducing WMDP scores.
D.2.5 FALCON
@@
For FALCON’s implementation, we maintained comparable learning rate ranges and number of iterations to RMU. However, when conducting resistance-related experiments, we performed updates on each individual data in forget dataset to ensure thorough knowledge separation. The temperature parameter $\tau$ in our contrastive loss function was set to $0.7$. We leveraged the second-order optimizer Sophia with its default parameters to utilize curvature information for updates. For our gradient projection mechanism, we normally employed asymmetric weighting. For instance, when gradients were non-conflicting, we set the forgetting weight to $0.8$ and retention weight to $1.2$; in cases of gradient conflict, these values were adjusted to $0.5$ and $1.5$ respectively. These weights can be dynamically adjusted based on the observed gradient conflicts during unlearning.
D.3 Entity and Copyrighted Content Unlearning
D.3.1 Open Unlearning Framework
The Open Unlearning Framework [16] provides a unified and extensible platform for evaluating machine unlearning methods in large language models. Developed by Locus Lab, it integrates both the TOFU and MUSE benchmarks, supporting experiments on synthetic and real-world datasets. The framework includes a range of unlearning algorithms and evaluation metrics, enabling researchers to systematically assess unlearning quality and model utility within a consistent environment. Our implementations of entity and copyrighted content unlearning are based on this Github [^4] https://github.com/locuslab/open-unlearning.
TOFU Benchmark
The Task of Fictitious Unlearning (TOFU) benchmark [56] is designed to evaluate the ability of large language models to selectively unlearn specific entity information while preserving overall model utility. TOFU introduces a synthetic dataset comprising biographies of fictitious authors, each containing detailed attributes such as birthplace, birth year, genre, and awards. During the unlearning experiments, a subset of authors (1%, 5%, or 10%) is designated as the Forget Set, while the rest form the Retain Set. To measure unlearning effectiveness, TOFU employs two main evaluation metrics. The Forget Quality is assessed using the Kolmogorov-Smirnov (KS) test, where a higher p-value indicates that the distribution of the unlearned model’s outputs becomes statistically closer to that of a model trained without the Forget Set. Model Utility evaluates how well the unlearned model retains knowledge about the Retain Set, real-world facts, and external author data. It is calculated as the harmonic mean of three performance indicators: answer probability, truth ratio, and ROUGE recall. This comprehensive design enables TOFU to rigorously evaluate the trade-offs between effective unlearning and model utility preservation under controlled experimental settings.
MUSE Benchmark
The Machine Unlearning Six-Way Evaluation (MUSE) benchmark [72] offers a comprehensive framework to assess machine unlearning in large language models, particularly focusing on real-world copyrighted and sensitive content. Unlike TOFU’s synthetic approach, MUSE targets naturally occurring datasets such as books and news articles, thus evaluating unlearning performance in more realistic and legally relevant scenarios. MUSE introduces several key evaluation dimensions. No Verbatim Memorization requires that the model does not reproduce exact text from the deleted data, preventing direct memorization. No Knowledge Memorization ensures that the model does not retain factual information derived solely from the forgotten data, even when rephrased. Utility Preservation emphasizes that the model should maintain its overall performance on unrelated tasks, ensuring that targeted unlearning does not degrade its general capabilities.
@@ E.2 Performance Breakdown Analysis of MMLU and WMDP
Humanities $6.85\pm 0.68 55$$5.86\pm 0.68$
   Formal Logic $5.24\pm 4.45 44$$4.44\pm 4.44$
@@
E.3 Computational Efficiency Comparison of FALCON and Other Baselines
The computational efficiency of different unlearning methods is assessed in Table 7, where we compare training runtime, processing throughput (samples/second), and optimization speed (steps/second) across all methods over 10 epochs on the TOFU benchmark. For fair comparison, all methods were implemented using first-order optimizers and evaluated under identical experimental conditions and framework (same hardware, batch sizes, and dataset configurations) [16]. FALCON achieves competitive efficiency with 13.94 seconds, processing 28.69 samples per second and completing 0.72 optimization steps per second in comparison to all baselines. These results confirm that our approach balances unlearning effectiveness with practical computational efficiency, making it suitable for real-world unlearning applications.
Table 7: Unlearning efficiency comparison of all unlearning methods over 10 epochs on TOFU.
Method Train Runtime (s) $\downarrow$ Samples/s $\uparrow$ Steps/s $\uparrow$
GA 8.71 45.94 1.15
GradDiff 19.65 20.36 0.51
NPO 30.83 12.97 0.32
IdkDPO 49.86 8.02 0.20
RMU 15.75 25.40 0.64
FALCON 13.94 28.69 0.72
E.4 Computational Efficiency of MI-guided parameter selection
We further analyze the computational efficiency of the proposed MI-guided parameter selection method on the TOFU benchmark using the Llama-3.2 backbone. The goal is to evaluate how MI estimation scales with different sample sizes while maintaining 95% PCA dimensionality retention for stable computation. We measured the runtime and identified the optimal intervention layer across different sample proportions (10%–100%). Results are summarized in Table 8 and 9, demonstrating consistent layer selection and manageable computational overhead.
Table 8: MI-guided method cost analysis on TOFU.
Sample Size Time (s) Optimal Layer
10% 67 3
30% 79 3
50% 110 3
70% 165 3
100% 260 3
Table 9: Normalized MI values across layers on full sample (lower indicates better disentanglement).
Layers 0–7
Layer 0 1 2 3 4 5 6 7
Normalized MI 0.78 0.43 0.19 0.00 0.18 0.07 0.05 0.14
Layers 8–15
Layer 8 9 10 11 12 13 14 15
Normalized MI 0.23 0.41 0.51 0.65 0.65 0.68 0.80 1.00
The results demonstrate that MI computation scales linearly with sample size while maintaining consistent optimal layer selection across data proportions, confirming the stability and efficiency of the MI-guided estimation. The identified optimal layer also aligns with the best empirical performance, indicating a strong correspondence between MI analysis and unlearning behavior. Overall, the proposed MI-guided mechanism offers a computationally efficient and scalable foundation for layer-level unlearning with stable and interpretable performance across dataset scales.
E.5 Ablation Study Analysis
To validate the effectiveness of FALCON’s components, we conduct ablation studies on Yi-6B-Chat. The baseline demonstrates a solid performance of 27.5% on WMDP and 60.3% on MMLU. Replacing the contrastive loss with RMU’s loss function (w/o Loss) renders unlearning ineffective, emphasizing the necessity of the contrastive mechanism for precise knowledge separation. While removing gradient projection (w/o GP) or replacing POVs with random vectors (w/o POVs) has a minor impact on unlearning but degrades knowledge retention and makes the model more vulnerable to Jailbreaking attacks [55], highlighting their critical role in preserving model utility and robustness. These results empirically confirm that each component is essential for FALCON’s success in achieving precise unlearning while maintaining general model performance.
Table 10: Impact of component omission on performance.
Variant Omit WMDP ($\downarrow$) MMLU ($\uparrow$)
Baseline 27.5 60.3
w/o Loss 50.7 61.4
w/o GP 27.4 58.4
w/o POVs 27.6 57.6
E.6 Evaluation of Recovery Resistance in Chat Settings
@@
E.7 Comparative Analysis of Knowledge Recovery Resistance
To further evaluate the robustness of FALCON against knowledge recovery attacks, we compare it with the representative unlearning baseline RMU. Unlike RMU, which relies on random perturbations in parameter space, FALCON introduces fine-grained representational shifts through contrastive orthogonal unalignment mechanism. This design steers activations away from dominant subspaces containing unwanted knowledge while preserving retention-sensitive directions, effectively creating geometric separation that renders forgotten representations structurally inaccessible. Table 12 summarizes the comparison on both the Bio and Cyber domains under identical recovery settings. The results demonstrate that FALCON maintains consistently lower recovery accuracy, indicating stronger resistance to reactivation of forgotten knowledge.
Table 12: Comparison of robustness performance under knowledge recovery attempts.
Method Original Unlearned Recovery
RMU (Bio) 65.4 50.8 58.5 $\pm$ 3.2
FALCON (Bio) 65.4 27.7 28.1 $\pm$ 0.5
RMU (Cyber) 42.6 33.5 41.8 $\pm$ 2.1
FALCON (Cyber) 42.6 25.3 25.5 $\pm$ 0.8
E.8 Example of Jailbreaking Prefixes Generated by Enhanced GCG